New Home, Inc. invested in the following bond issued by Old Home, Inc. on January 1, Yr1. The company’s amortization schedule appears below:

Date

Cash Flow

Interest

Revenue

Amortized Discount

Bond

Discount

Book Value

of Bond

Fair Value

1/1/Yr1

$16,158

?

6/30/Yr1

?

$11,692

$1,692

$14,466

$235,534

$240,000

12/31/Yr1

?

$11,777

$1,777

$12,689

$237,311

$242,000

6/30/Yr2

?

$11,866

$1,866

$10,824

$239,176

$244,000

12/31/Yr2

?

$11,959

$1,959

$ 8,865

$241,135

$246,000

6/30/Yr3

?

$12,057

$2,057

$ 6,808

$243,192

$248,000

12/31/Yr3

?

$12,160

$2,160

$ 4,649

$245,351

$250,000

6/30/Yr4

?

$12,268

$2,268

$ 2,381

$247,619

$252,000

12/31/Yr4

?

$12,381

$2,381

$         0

?

$254,000

Use the amortization table to answer the following questions.

a) What is the maturity value of the bond?

b) How much cash did New Home, Inc. pay upon purchasing the bond?

c) If you were an investor that had purchased one fourth of the notes, how much interest would you receive during Yr3? How much interest would you have earned?

d) What interest rate and how many periods were used to determine the present value of the bond.

e) If New Home sold this bond (classified as trading security) on 1/10/Yr3 for $246,500, prepare the necessary journal entries for the sale.

f) If New Home sold this bond (classified as available for sale security) on 1/10/Yr3 for $244,500, prepare the necessary journal entries for the sale.
Answer #1

a) . Maturity value of the bond will be equal to face value of the bond and face value of the bond = book value of bond + discount in bond

This, face value = $235534+$14466 = $250000.

b) . New home inc would have paid the face value less discount i.e. book value of bond. Thus, amount paid by new home inc. in purchasing the bond = $250000-$16158 = $233842.

c). Note that interest revenue given is not the coupon amount received on bond since the interest revenue given is consistently changing every period whereas coupon amount remain constant. Thus, we have coupon amount = $11692-$1692 = $10000 , thus, this interest revenue is the amount of interest earned .

So for y3 amount of interest earned = (12057+12160)/4 = $6054.25

Interest received = (10000+10000)/4 = $5000.

d) . Period used = 2x4 = 8

Whereas coupon rate = (10000/250000)X100X2 = 8%.

e ) . If new home sold this bond for $246500 in 1/10/y3 and it is held aa trading security then , note that we will record recognized gain or loss based on proceeds from sale - value recorded at the end of immediately preceding balance sheet date i.e. 12/31/y2.

Thus, following will be journal entry

Date general journal Debit credit
1/10/yr 3

cash

Investment (trading security)

Recognized gain on sale of investment

(To sale of trading securities and recognized gain recorded as 246500+246000)

246500

246000

500

Note :- last updated price of bond will be aa as per fair value aa on 12/31/yr 2 i.e. $246000.

f) . In case security is held aa avable for sale any change in value on Balance sheet is passed through equity and not through income statement, and finally when such security is of then any gain or loss must be recorded in income statement and any gain or loss previously recorded through equity shall be cancelled. Following will be the journal entry :-

Date general journal debit credit
1/10/yr3

Cash

Unrecognized gain on bond (equity)[see note 2)

Investment

Gain on sale of bond (investment)

(To sale of bond held as available for sale)

244500

12158

246000

10658

Note

1 Original gain on sale = 244500-233842(amount initially paid on acquiring bond)

2 Unrecognized gain on sale of bond (equity) to be cancelled = 246000-233842 = $12158